OpenStudy (anonymous):

A bag contains 4 red balls and 8 yellow balls. Five are selected one at a time at random without replacement. Find P( 4/5 are red)

OpenStudy (anonymous):

P(4/5 are red)= 4/12 * 3/11 * 2/10 * 1/9 * 8/8 ( 4 red and 1 yellow) = 0.002

OpenStudy (anonymous):

yeah, i did that, but the answer is 0.5, idk

OpenStudy (anonymous):

i mean i got 0.002, but the real answer is 0.5

OpenStudy (anonymous):

Hmm i dont see how it can be 0.5 maybe the answer given is wrong, or we are wrong. i honestly have no idea.